OpenStudy (anonymous):

Find S20 for the arithmetic series 8 + 16 + 24 + 32 +… super confused cause my options for answers are 1)1672 2)1688 3)1664 4)1680 shouldn't the answer be 160?

OpenStudy (anonymous):

First you have to find the 20th term... an = a1 + (n - 1)d n = 20 a1 = 8 d = 8 now lets sub... a20 = 8 + (20 - 1) * 8 a20 = 8 + 19 * 8 a20 = 8 + 152 a20 = 160 s20 is the sum of all the number till 20. so we will use the sum formula... sn = n(a1 + an) / 2 s20 = 20 (8 + 160) / 2 s20 = 20 (168) / 2 s20 = 3360/2 s20 = 1680 any questions ?